Describe the complement of the given event.

73% of nineteen-year-old males are at least 169 pounds.
Answer #1

A complement event of an event X is the set of all events which are excluded from event X. Therefore, in this case,

P(A) + P(Ac) = 1

P(Ac) = 1 - P(A), where Ac: Complementary event to event A

Complementary event: 27% of nineteen-year-old males are less than 169 pounds [P(Ac) = 1 - 0.73 = 0.27 = 27%].